# Weakly Supervised Gaussian Contrastive Grounding with Large Multimodal Models for Video Question Answering

### v3

@@ Abstract.
Video Question Answering (VideoQA) aims to answer natural language questions based on the information observed in videos. Despite the recent success of Large Multimodal Models (LMMs) in image-language understanding and reasoning, they deal with VideoQA insufficiently, by simply taking uniformly sampled frames as visual inputs, which ignores question-relevant visual clues. Moreover, there are no human annotations for question-critical timestamps in existing VideoQA datasets. In light of this, we propose a novel weakly supervised framework to enforce the LMMs to reason out the answers with question-critical moments as visual inputs. Specifically, we first fuse the question and answer pairs as event descriptions to find multiple keyframes as target moments and pseudo-labels, with the visual-language alignment capability of the CLIP models. With these pseudo-labeled keyframes as additionally weak supervision, we devise a lightweight Gaussian-based Contrastive Grounding (GCG) module. GCG learns multiple Gaussian masks to characterize the temporal structure of the video, and sample question-critical frames as positive moments to be the visual inputs of LMMs. Extensive experiments on several benchmarks verify the effectiveness of our framework, and we achieve substantial improvements compared to previous state-of-the-art methods.
@@
[…]
Figure 1. We argue that the information in uniformly sampled frames is insufficient for LMMs to answer the question correctly. Therefore, we utilize the fused event description to provide additional weak supervision and generate weight distributions for each video moment. We align the positive description-moment pairs while pushing away negative ones. In this way, we can effectively select question-critical moments for LMMs to reason out the answer.
@@ 1. Introduction
To summarize, we make the following contributions: (1) We propose a weakly supervised grounding framework for VideoQA, by utilizing the alignment capability of the CLIP models to provide pseudo-labeled timestamps of keyframes without human-labor annotated costs. (2) We devise the Gaussian-based Contrastive Grounding (GCG) for weakly-grounded selection of question-critical moments, enhancing the effectiveness and interpretability of LMMs when applied to VideoQA, by revealing which visual scenes result in the predicted answers. (3) We conduct extensive experiments to verify the effectiveness of our proposed method, and achieve substantial improvements on six challenging VideoQA benchmarks including NExT-QA, Causal-VidQA, Intent-QA, ActivityNet-QA, MSVD-QA, and MSRVTT-QA.
@@ 2.2. Temporal Grounding in VideoQA
Early VideoQA benchmarks (Yu et al., 2019; Xu et al., 2017; Jang et al., 2017) focus on descriptive questions (e.g., [what’s the man doing]) within short video clips, rarely going beyond a recognition of the objects and actions. Instead, more recent VideoQA benchmarks (Li et al., 2022b; Xiao et al., 2021; Li et al., 2023d) like NExT-QA (Xiao et al., 2021) emphasize counterfactual, temporal, and causal reasoning involving multiple entities and relations, demanding the ability to uncover the causes behind specific events within longer videos, necessitating localizing a text query to specific moments.
In light of this, ATP (Buch et al., 2022) utilizes the tool of atemporal probe to select a single frame without temporal information for downstream tasks. MIST (Gao et al., 2023) and TranSTR (Li et al., 2023e) fuse frames with the mechanism of adaptive temporal rationalization and iterative spatial-temporal attention, respectively. NExT-GQA (Xiao et al., 2023) constructs grounding labels in the test set of NExT-QA and uses a single Gaussian mask to fuse frames along the temporal dimension. SeViLA (Yu et al., 2023), similar to us, utilizes the LMM (BLIP-2) for VideoQA. However, SeViLA uses two LMMs to generate pseudo-labels and answer questions respectively, with extra pre-training on TSG datasets (Lei et al., 2021) and a multi-stage training scheme. Different from previous works, we utilize the CLIP (Radford et al., 2021; Fang et al., 2023) models to automatically provide weak supervision for grounding, and our lightweight GCG module learns multiple Gaussian masks to generate both positive and negative moments in an end-to-end manner, with an additionally contrastive objective to distinguish positive ones from negative ones for frame selection.
@@ 3. Preliminary: LMMs for VideoQA
We take InstructBLIP (Dai et al., 2023) as an example to illustrate how LMMs deal with VideoQA. Similar to other LMMs, InstructBLIP approaches VideoQA as a text generation task conditioned on the question $\mathcal{Q}$ and video $\mathcal{V}$ with $T$ frames, and predicts the answer $\mathcal{A}$ by the following three steps:
(1) The Vision Transformer (Dosovitskiy et al., 2021) in EVA-CLIP (Fang et al., 2023) serves as the frozen image encoder to extract embeddings of each frame individually, and obtains $\textbf{E}=\left\{\textbf{e}_{1},\textbf{e}_{2},\cdots,\textbf{e}_{T}\right\},\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}},\textbf{e}_{t}\in\mathbb{R}^{N_{I}\times D_{I}}$, where $t$ denotes the $t$-th frame, $N_{I}$ is the patch number of each frame (including the class token), and $D_{I}$ is the embedding dimension. To mitigate computational costs, existing LMMs uniformly sample $K$ frames ($K\ll T$) to represent the video, resulting in the sampled $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$.
(2) A trainable Q-former serves as the connection module to bridge the modality gap between vision and language. It takes frame embeddings $\hat{\textbf{E}}$ as inputs and outputs a set of fixed-length frame tokens $\textbf{F}=\left\{\textbf{f}_{1},\textbf{f}_{2},\cdots,\textbf{f}_{K}\right\},\textbf{F}\in\mathbb{R}^{K\times N_{C}\times D_{C}},\textbf{f}_{t}\in\mathbb{R}^{N_{C}\times D_{C}}$, where $N_{C}$ is the token number of each frame ($N_{C}\ll N_{I}$, e.g., $N_{C}=32$ and $N_{I}=257$ in InstructBLIP), and $D_{C}$ is the dimension of the connection module.
@@ 4.2. Pseudo Labels for Temporal Grounding
Considering the powerful visual-language alignment ability of EVA-CLIP, as in Figure 2 (b), we utilize its joint-trained image and text encoder to provide pseudo labels for timestamps of question-critical moments as weak supervision.
@@
Pseudo Labels Generation. As in Figure 2 (b), we represent the video and event description with the class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ and $\textbf{d}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ respectively, to obtain the weakly labeled question-critical timestamps. In detail, we compute the cosine similarities between $\textbf{E}_{[CLS]}$ and $\textbf{d}_{[CLS]}$ and get the similaruity scores $s\in\mathbb{R}^{T}$, which recording the relevance between each frame and the event description. Then, we choose the indexes of the highest Top-$K$ scores in $s$ as $w\in\mathbb{R}^{K}$, where each element $w_{k}\in\left\{1,2,\cdots,T\right\}$, to be the timestamps of question-critical frames. We verify the effectiveness of the weakly grounded $w$ in the analysis 5.3.
@@ 4.3. Gaussian Generator
Recent research (Xiao et al., 2023) highlights the superiority of end-to-end Gaussian mask learning for grounding in VideoQA. Motivated by this, we design the effective weakly supervised Gaussian-based Contrastive Grounding (GCG). However, unlike (Xiao et al., 2023) to generate a single Gaussian mask, we generate multiple Gaussian masks to characterize the multi-event temporal structure of the video. Moreover, our GCG is optimized from both the QA supervision $\mathcal{L}_{vqa}$ and weakly labeled supervision $\mathcal{L}_{reg}$ and $\mathcal{L}_{con}$, which is more effective for discovering the question-critical moments.
@@ 4.4. Contrastive Grounding
$\mathrm{SUM}=\sum_{i=1}^{N_{intra}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{intra}_{i}/\tau)\ +$
$\sum_{j=1}^{N_{inter}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{inter}_{j}/\tau)$
@@ 5.1. Datasets
NExT-QA (Xiao et al., 2021) contains 5.4k videos with an average length of 44s and 52k QA pairs, including question types of description, causal, and temporal. Intent-QA (Li et al., 2023d) focuses on intent reasoning in daily social activities, with more than 4.3k videos and 16k QA pairs, including question types of causal-why, causal-how, and temporal. Causal-VidQA (Li et al., 2022b) selects 27k video clips and asks 108k questions, including types of description, explanation, prediction, and counterfactual. MSVD-QA and MSRVTT-QA (Xu et al., 2017) emphasize the description of video objects, activities, and their attributes, with 50k QA pairs over 1,970 videos and 243K QA pairs over 10K videos respectively. ActivityNet-QA (Yu et al., 2019) consists of 58k QA pairs on 5.8k long web videos, with an average length of 180 seconds. NExT-QA, Intent-QA, and Causal-VidQA use a multi-choice setting to test temporal reasoning with causal and commonsense relations. MSVD-QA, MSRVTT-QA, and ActivityNet-QA employ an open-ended setting, focusing on descriptive questions of different elements.
@@ 5.2. Implementation Details
We choose InstructBLIP (Dai et al., 2023) and BLIP-2 (Li et al., 2023b) as our LMM for their representative structure and widespread use, with EVA-CLIP (Fang et al., 2023) as the image encoder, Q-former (Li et al., 2023b) as the connection module, and Vicuna (Zheng et al., 2023) or FLAN-T5 (Chung et al., 2022) as the large language model. Following previous works (Yu et al., 2023; Li et al., 2023e, b; Dai et al., 2023), we sample each video as a sequence of $T=32$ frames and select $K=4$ frames as visual inputs. The number of negative samples is $N_{intra}=16$ and $N_{inter}=32$. The number of transformer encoder layers in the Gaussian generator is 2, with the hidden size $D_{G}=256$. For the hyperparameters, we set $\sigma=0.2$, $\tau=0.1$, $\alpha_{1}=\alpha_{2}=0.1$. During training, we keep the parameters of the image encoder, LLM, and text encoder frozen. We use AdamW to optimize the model with a learning rate of $1e^{-5}$ and the strategy of mixed precision. For multi-choice datasets like NExT-QA, Causal-VidQA, and Intent-QA, we use BLIP2-FLAN-T5-XL and IntructBLIP-FLAN-T5-XL. For open-ended datasets like MSVD-QA, MSRVTT-QA, and ActivityNet-QA, we use IntructBLIP-Vicuna-7B.
@@
5.4. Main Results
Table 1 and 2 show the superiority of our method. For the multi-choice setting, we achieve an accuracy of 74.6%, 72.1%, and 73.1% in NExT-QA, Causal-VidQA, and Intent-QA respectively. For the open-ended setting, we achieve an accuracy of 61.7%, 49.5%, and 49.9% in MSVD-QA, MSRVTT-QA, and ActivityNet-QA respectively. To ensure a fair comparison, we also apply the same settings to get the results for the vanilla InstructBLIP on these datasets as baselines, with $K=4$ frames uniformly sampled from the $T=32$ frames as visual inputs.
@@ 5.4. Main Results
Although the baseline InstructBLIP performed fair on these datasets, our proposed GCG showed a significant improvement, particularly in questions that require complex causal-temporal reasoning (+2.1% and +2.7% for Tem and Cau in NExT-QA, +2.3% for Cou in Causal-VidQA). When both using BLIP-2 as the baseline model, GCG still performs better than SeViLA (Yu et al., 2023) (73.6% vs 71.5% in NExT-QA) without extra pre-training or a multi-stage training scheme. Moreover, despite not having undergone video-text pretraining, our method still surpasses those large-scale pre-trained models (e.g., HiTea, COSA, VALOR) on MSVD-QA, MSRVTT-QA, which primarily feature straightforward questions and short videos (10-15s). We also observe that the improvements on ActivityNet-QA (+3.6%) are much larger than the MSVD-QA (+2%) and MSRVTT-QA (+2.5%). This can be attributed to the average video length of ActivityNet-QA being 180 seconds, which is much longer than MSVD-QA (10s) and MSRVTT-QA (15s), emphasizing the necessity of discovering question-critical moments more with our method.
@@ 5.5. Ablation Studies
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ and $\mathcal{L}_{vqa}+\mathcal{L}_{con}$ match equally in accuracy that consistently surpasses baseline and $\mathcal{L}_{vqa}$. $\mathcal{L}_{reg}$ is responsible for regularizing the indexes of peaks in $p$ to approximate the timestamps of weakly grounded $w$,
    Figure 6. Ablation studies on the number of negative samples.
@@
    and $\mathcal{L}_{con}$ ensures the maximization between the selected moments and event descriptions.
Number of frames $T$ and Gaussian widths $\sigma$. We evaluate how the number of overall sampled frames $T$ and the widths $\sigma$ of Gaussian functions will influence the performance. Figure 5 (a) indicates that performance improves as more frames are included, however, beyond a certain threshold ($T=48$), there is a performance drop. This suggests that too many frames may introduce redundancy and noise, while too few frames miss important information. As for $\sigma$ in Equation 4, it essentially decides the width and the degree of dispersion of the Gaussian distribution $g$. A larger $\sigma$ generates more dispersed $g$ with a more exploratory $p$ and vice versa. In Figure 5 (b), we vary $\sigma$ from 0.1 to 0.7 and observe that the performance fluctuates in a range of [73.9, 74.6] for NExT-QA and [61.4, 61.7] for MSVD-QA. The performance of NExT-QA is more sensitive to $\sigma$, and we argue that this is because the average video length of NExT-QA (44s) is much longer than MSVD-QA (10s).
@@
Ablations on negative sample numbers $N_{intra}$ and $N_{inter}$. We also investigate how the number of intra-negative samples $N_{intra}$ and inter-negative samples $N_{inter}$ influence the performance. As verified in Figure 6 (a), sampling enough intra-negative moments is beneficial to mining the positive moments for reasoning. Figure 6 (b) shows that the performance gains also increase as the number of inter-video negative moments increases. Moreover, the impact of intra-negative moments within the same video is larger than inter-negative moments, because the scenes in the same video are more confusing to distinguish the true question-critical moments. We also observe that the performance decreases after the $N_{intra}$ and $N_{inter}$ reach a specific value of 16 and 32 respectively. We argue that selecting excessive negative moments tends to distract positive moments and therefore degrades the performance. We finally set the $N_{intra}=16$ and $N_{inter}=32$ which balance well the performance and computational cost.
Ablations on hidden size $D_{R}$ and layer number $N$. To determine the optimal configuration for our Gaussian generator, we set different layer number $N$ and hidden size $D_{R}$ for the transformer encoder in the Gaussian generator. The results of these variants on NExT-QA are shown in Table 4. We can see that when the layer number of the encoder is fixed to $N=2$, the model achieves the best performance of $74.8\%$ with $D_{G}=1024$. However, such improvement in performance comes at the cost of a significant increase in the number of parameters (27.6M). To ensure the flexibility and adaptability of our method, we choose $D_{G}=256$ as our default setting, which achieves a balance between fair performance (74.6%) and much fewer parameter quantities (2.0M). We can also observe that when the hidden size is fixed to $D_{G}=256$ and the layer number $N\geq 2$, the change in $N$ has a relatively minor impact on performance. For better computation efficiency, we adopt $N=2$ for our Gaussian generator.
Further LoRA tuning to push better results. We note that LMMs with GCG can be considered a strong model for VideoQA. To achieve better performance, we further add per-task LoRA tuning (Hu et al., 2021) for the frozen language model in LMMs (using a rank of 16), yielding new SOTA results on these VideoQA benchmarks. These new SOTA results in Table 5 indicate the extensibility and flexibility of the design of our GCG, which can be easily combined with other components for LMMs.

### v4

@@ Abstract.
Video Question Answering (VideoQA) aims to answer natural language questions based on the information observed in videos. Despite the recent success of Large Multimodal Models (LMMs) in image-language understanding, they deal with VideoQA insufficiently, by simply taking uniformly sampled frames as visual inputs, which ignores question-relevant visual clues. Moreover, there are no human annotations for question-critical timestamps in existing VideoQA datasets. In light of this, we propose a weakly supervised framework to enforce the LMMs to reason out the answers with question-critical moments as visual inputs. Specifically, we fuse the question and answer pairs as event descriptions to find multiple keyframes as target moments and pseudo-labels. With these pseudo-labeled keyframes as additionally weak supervision, we devise a lightweight Gaussian-based Contrastive Grounding (GCG) module. GCG learns multiple Gaussian masks to characterize the temporal structure of the video, and sample question-critical frames as positive moments to be the visual inputs of LMMs. Extensive experiments on several benchmarks verify the effectiveness of our framework, and we achieve substantial improvements compared to previous state-of-the-art methods. Codes will be available at link.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680826isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Visual content-based indexing and retrieval
Figure 1. The information in uniformly sampled frames is insufficient to answer the question correctly. We utilize the fused event description to provide weak supervision and generate weight distributions for each video moment. We align the positive description-moment pairs while pushing away negative ones.
@@ 1. Introduction
To summarize, we make the following contributions: (1) We propose a weakly supervised grounding framework for VideoQA, by utilizing the CLIP models to provide pseudo-labeled timestamps of keyframes without human-labor annotated costs. (2) We devise the Gaussian-based Contrastive Grounding (GCG) for weakly-grounded selection of question-critical moments, enhancing the effectiveness and interpretability of LMMs when applied to VideoQA, by revealing which visual scenes result in the predicted answers. (3) We conduct extensive experiments to verify the effectiveness of our proposed method, and achieve substantial improvements on six challenging VideoQA benchmarks including NExT-QA, Causal-VidQA, Intent-QA, ActivityNet-QA, MSVD-QA, and MSRVTT-QA.
@@ 2.2. Temporal Grounding in VideoQA
Early VideoQA benchmarks (Yu et al., 2019; Xu et al., 2017; Jang et al., 2017) focus on descriptive questions (e.g., [what’s the man doing]) within short video clips, rarely going beyond a recognition of the objects and actions. Instead, more recent VideoQA benchmarks (Li et al., 2022b; Xiao et al., 2021; Li et al., 2023d) like NExT-QA (Xiao et al., 2021) emphasize counterfactual, temporal, and causal reasoning involving multiple entities and relations, demanding the ability to uncover the causes behind specific events within longer videos, necessitating localizing a text query to specific moments. In light of this, ATP (Buch et al., 2022) utilizes the tool of atemporal probe to select a single frame without temporal information for downstream tasks. MIST (Gao et al., 2023) and TranSTR (Li et al., 2023e) fuse frames with the mechanism of adaptive temporal rationalization and iterative spatial-temporal attention, respectively. NExT-GQA (Xiao et al., 2023) constructs grounding labels in the test set of NExT-QA and uses a single Gaussian mask to fuse frames along the temporal dimension. SeViLA (Yu et al., 2023), similar to us, utilizes the LMM (BLIP-2) for VideoQA. However, SeViLA uses two LMMs to generate pseudo-labels and answer questions respectively, with extra pre-training on TSG datasets (Lei et al., 2021) and a multi-stage training scheme. Different from previous works, we utilize the CLIP (Radford et al., 2021; Fang et al., 2023) models to automatically provide weak supervision for grounding, and our lightweight GCG module learns multiple Gaussian masks to generate both positive and negative moments in an end-to-end manner, with an additionally contrastive objective to distinguish positive ones from negative ones for frame selection.
@@ 3. Preliminary: LMMs for VideoQA
We take InstructBLIP (Dai et al., 2023) as an example to illustrate how LMMs deal with VideoQA. InstructBLIP approaches VideoQA as a text generation task conditioned on the question $\mathcal{Q}$ and video $\mathcal{V}$ with $T$ frames, and predicts the answer $\mathcal{A}$ by the following three steps:
(1) The ViT (Dosovitskiy et al., 2021) in EVA-CLIP (Fang et al., 2023) serves as the frozen image encoder to extract embeddings of each frame individually, and obtains $\textbf{E}=\left\{\textbf{e}_{1},\textbf{e}_{2},\cdots,\textbf{e}_{T}\right\},\textbf{E}\in\mathbb{R}^{T\times N_{I}\times D_{I}},\textbf{e}_{t}\in\mathbb{R}^{N_{I}\times D_{I}}$, where $t$ denotes the $t$-th frame, $N_{I}$ is the patch number of each frame (including the class token), and $D_{I}$ is the embedding dimension. To reduce tokens, existing LMMs uniformly sample $K$ frames ($K\ll T$) to represent the video, resulting in the sampled $\hat{\textbf{E}}\in\mathbb{R}^{K\times N_{I}\times D_{I}}$.
(2) A trainable Q-former serves as the connection module to bridge the modality gap. It takes frame embeddings $\hat{\textbf{E}}$ as inputs and outputs a set of fixed-length frame tokens $\textbf{F}=\left\{\textbf{f}_{1},\textbf{f}_{2},\cdots,\textbf{f}_{K}\right\},\textbf{F}\in\mathbb{R}^{K\times N_{C}\times D_{C}},\textbf{f}_{t}\in\mathbb{R}^{N_{C}\times D_{C}}$, where $N_{C}$ is the token number of each frame ($N_{C}\ll N_{I}$, e.g., $N_{C}=32$ and $N_{I}=257$ in InstructBLIP), and $D_{C}$ is the dimension of the connection module.
@@ 4.2. Pseudo Labels for Temporal Grounding
Considering the powerful alignment ability of EVA-CLIP, as in Figure 2 (b), we utilize its joint-trained image and text encoder to provide pseudo labels for timestamps of question-critical moments.
@@
Pseudo Labels Generation. As in Figure 2 (b), we represent the video and event description with the class tokens $\textbf{E}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ and $\textbf{d}_{[CLS]}\in\mathbb{R}^{T\times D_{I}}$ respectively, to obtain the weakly labeled question-critical timestamps. In detail, we compute the cosine similarities between $\textbf{E}_{[CLS]}$ and $\textbf{d}_{[CLS]}$ and get the similaruity scores $s\in\mathbb{R}^{T}$, which recording the relevance between each frame and the event description. Then, we choose the indexes of the highest Top-$K$ scores in $s$ as $w\in\mathbb{R}^{K}$, where each element $w_{k}\in\left\{1,2,\cdots,T\right\}$, to be the timestamps of question-critical frames.
@@ 4.3. Gaussian Generator
Recent research (Xiao et al., 2023) highlights the superiority of Gaussian mask learning for grounding. Motivated by this, we design the weakly supervised Gaussian-based Contrastive Grounding (GCG). However, unlike (Xiao et al., 2023) to generate a single Gaussian mask, we generate multiple Gaussian masks to characterize the multi-event temporal structure of the video. Moreover, our GCG is optimized from both the QA supervision $\mathcal{L}_{vqa}$ and weak supervision $\mathcal{L}_{reg}$ and $\mathcal{L}_{con}$.
@@ 4.4. Contrastive Grounding
$\mathrm{SUM}=\sum_{i=1}^{N_{intra}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{intra}_{i}/\tau)\ +\sum_{j=1}^{N_{inter}}\mathrm{exp}(\textbf{d}_{[CLS]}\otimes\textbf{E}^{inter}_{j}/\tau)$
@@ 5.1. Datasets
NExT-QA (Xiao et al., 2021) contains 5.4k videos with an average length of 44s and 52k QA pairs, including question types of description, causal, and temporal. Intent-QA (Li et al., 2023d) focuses on intent reasoning in daily social activities, with more than 4.3k videos and 16k QA pairs, including question types of causal-why, causal-how, and temporal. Causal-VidQA (Li et al., 2022b) selects 27k video clips and asks 108k questions, including types of description, explanation, prediction, and counterfactual. MSVD-QA and MSRVTT-QA (Xu et al., 2017) emphasize the description of video objects, activities, and their attributes, with 50k QA pairs over 1,970 videos and 243K QA pairs over 10K videos respectively. ActivityNet-QA (Yu et al., 2019) consists of 58k QA pairs on 5.8k long web videos, with an average length of 180 seconds.
@@ 5.2. Implementation Details
We choose InstructBLIP (Dai et al., 2023) and BLIP-2 (Li et al., 2023b) as our LMM for their representative structure and widespread use, with EVA-CLIP (Fang et al., 2023) as the image encoder, Q-former (Li et al., 2023b) as the connection module, and Vicuna (Zheng et al., 2023) or FLAN-T5 (Chung et al., 2022) as the large language model. Following previous works (Yu et al., 2023; Li et al., 2023e, b; Dai et al., 2023), we sample each video as a sequence of $T=32$ frames and select $K=4$ frames as visual inputs. The number of negative samples is $N_{intra}=16$ and $N_{inter}=32$. The number of transformer encoder layers in the Gaussian generator is 2, with the hidden size $D_{G}=256$. For the hyperparameters, we set $\sigma=0.2$, $\tau=0.1$, $\alpha_{1}=\alpha_{2}=0.1$. During training, we keep the parameters of the image encoder, LLM, and text encoder frozen. We use AdamW to optimize the model with a learning rate of $1e^{-5}$ and the strategy of mixed precision.
@@
5.4. Main Results
Table 1 and 2 show that, for the multi-choice setting, we achieve an accuracy of 74.6%, 72.1%, and 73.1% in NExT-QA, Causal-VidQA, and Intent-QA respectively. For the open-ended setting, we achieve an accuracy of 61.7%, 49.5%, and 49.9% in MSVD-QA, MSRVTT-QA, and ActivityNet-QA respectively. To ensure a fair comparison, we also apply the same settings to get the results for the vanilla InstructBLIP on these datasets as baselines, with $K=4$ frames uniformly sampled from the $T=32$ frames as visual inputs.
Although the baseline InstructBLIP performed fair on these datasets, our proposed GCG showed a significant improvement, particularly in questions that require complex causal-temporal reasoning (+2.1% and +2.7% for Tem and Cau in NExT-QA, +2.3% for Cou in Causal-VidQA). When both using BLIP-2 as the baseline model, GCG still performs better than SeViLA (Yu et al., 2023) (73.6% vs 71.5% in NExT-QA) without extra pre-training or a multi-stage training scheme.
@@ 5.4. Main Results
Figure 6. Ablation studies on the number of negative samples.
Moreover, despite not having undergone video-text pretraining, our method still surpasses those large-scale pre-trained models (e.g., HiTea, COSA, VALOR) on MSVD-QA, MSRVTT-QA, which primarily feature straightforward questions and short videos (10-15s). We also observe that the improvements on ActivityNet-QA (+3.6%) are much larger than the MSVD-QA (+2%) and MSRVTT-QA (+2.5%). This can be attributed to the average video length of ActivityNet-QA being 180 seconds, which is much longer than MSVD-QA (10s) and MSRVTT-QA (15s), emphasizing the necessity of discovering question-critical moments more with our method.
@@ 5.5. Ablation Studies
  • $\mathcal{L}_{vqa}+\mathcal{L}_{reg}$ and $\mathcal{L}_{vqa}+\mathcal{L}_{con}$ match equally that surpasses baseline and $\mathcal{L}_{vqa}$. $\mathcal{L}_{reg}$ is responsible for regularizing the indexes of peaks in $p$ to approximate the timestamps of weakly grounded $w$, and $\mathcal{L}_{con}$ ensures the maximization between the selected moments and event descriptions.
@@
Number of frames $T$ and Gaussian widths $\sigma$. Figure 5 (a) indicates that performance improves as more frames are included, however, beyond a certain threshold ($T=48$), there is a performance drop. This suggests that too many frames may introduce redundancy and noise, while too few frames miss important information. As for $\sigma$ in Equation 4, it essentially decides the width and the degree of dispersion of the Gaussian distribution $g$. A larger $\sigma$ generates more dispersed $g$ with a more exploratory $p$ and vice versa. In Figure 5 (b), we vary $\sigma$ from 0.1 to 0.7 and observe that the performance fluctuates in a range of [73.9, 74.6] for NExT-QA and [61.4, 61.7] for MSVD-QA. The performance of NExT-QA is more sensitive to $\sigma$, and we argue that this is because the average video length of NExT-QA (44s) is much longer than MSVD-QA (10s).
@@
Ablations on negative sample numbers $N_{intra}$ and $N_{inter}$. As verified in Figure 6 (a), sampling enough intra-negative moments is beneficial to mining the positive moments for reasoning. Figure 6 (b) shows that the performance gains also increase as the number of inter-video negative moments increases. Moreover, the impact of intra-negative moments within the same video is larger than inter-negative moments, because the scenes in the same video are more confusing to distinguish the true question-critical moments. We also observe that the performance decreases after the $N_{intra}$ and $N_{inter}$ reach a specific value of 16 and 32 respectively. We argue that selecting excessive negative moments tends to distract positive moments and therefore degrades the performance.
Ablations on hidden size $D_{R}$ and layer number $N$. To determine the optimal configuration for our Gaussian generator, we set different layer number $N$ and hidden size $D_{R}$ for the transformer encoder in the Gaussian generator. The results of these variants on NExT-QA are shown in Table 4. We can see that when the layer number of the encoder is fixed to $N=2$, the model achieves the best performance of $74.8\%$ with $D_{G}=1024$. However, such improvement in performance comes at the cost of a significant increase in the number of parameters (27.6M). To ensure the flexibility and adaptability of our method, we choose $D_{G}=256$ as our default setting, which achieves a balance between fair performance (74.6%) and much fewer parameter quantities (2.0M). We can also observe that when the hidden size is fixed to $D_{G}=256$ and the layer number $N\geq 2$, the change in $N$ has a relatively minor impact on performance. For better computation efficiency, we adopt $N=2$.
Further LoRA tuning to push better results. We note that LMMs with GCG can be considered a strong model for VideoQA. To achieve better performance, we further add LoRA tuning (Hu et al., 2021) for the frozen language model in LMMs (using a rank of 16), yielding new SOTA results on these VideoQA benchmarks. Results in Table 5 indicate the extensibility and flexibility of the design of our GCG, which can be easily combined with other components for LMMs.